OpenStudy (anonymous):

I need help! simplifying 12x-96/x^2-4x-32 /6x^2-6x-12/3x+3

OpenStudy (sriram):

plz use the option of equation for long equations

OpenStudy (anonymous):

\[(12x-9)6/(x^2-4x) - 32/(6x^2-6x) -12/(3x+3) ?\]

OpenStudy (anonymous):

error (12x - 96)

OpenStudy (anonymous):

12x-96 over...x^2-4x-32 ..over 6x^2-6x-12/3x+3

OpenStudy (anonymous):

I guess you can solve it like that?

OpenStudy (anonymous):

is my suggestion correct?

OpenStudy (anonymous):

my options for answeres are...6(x+4)/x-2 orrr.......x+4/6(x-2)....or 6/(x+4)(x-2)....or (x+4)(x-2)/6

OpenStudy (anonymous):

15x^2-4x^3-4x-272/3 if your question at first is right

OpenStudy (anonymous):

thats not even a option

OpenStudy (anonymous):

i dont think my suggestion is correct - the algebra is horrendous and none of those options would fit

OpenStudy (phi):

\[\frac{12x-96}{x ^{2}-4x-32} \frac{3x+3}{6x^{2}-6x-12}\] Is this it?? I re-wrote it a bit using \[\frac{a/b}{c/d}=\frac{a}{b}\frac{d}{c}\] (Multiply top and bottom by d/c)

OpenStudy (phi):

To simplify, you have to start factoring...